A 2-kg disk has a radius of 18 cm and rotates with an angular acceleration of 12 rad/s2

about

an axis through its center and perpendicular to the plane of the disk. Determine the tangential

force at the rim of the disk


1
Expert's answer
2021-03-14T19:18:48-0400

By the definition of the torque, we have:


"\tau=FR,"

here, "F" is the tangential force at the rim of the disk, "r" is the radius of the disk.

From the other hand,


"\tau=I\alpha."

The moment of inertia of the disk can be written as follows:


"I=\dfrac{1}{2}MR^2."

Finally, we have:


"FR=\dfrac{1}{2}MR^2\alpha,""F=\dfrac{1}{2}MR\alpha,""F=\dfrac{1}{2}\cdot2\ kg\cdot0.18\ m\cdot12\ \dfrac{rad}{s^2}=2.16\ N."